Ann invested her savings in two investment funds. The 6000 that she invested in Fund A returned a 5% profit. The amount that she invested in Fund B returned a 1% profit. How much did she invest in Fund B, if both funds together returned a 4% profit?
Answer by mananth(16946)

percent ---------------- Amount $
Fund A 5 ---------------- 6000
Fund B 1 ---------------- x
Total 4 ---------------- 6000 + x

5*6000 +1x=4(6000+x)
30000 + 1x = 24000 +4x
x-4x=-30000+24000
-3x=-6000
/-3
x=2000
$2000 in Fund B
m.ananth@hotmail.ca
Check
2000@1%=20
6000@6%= 300
Total 320
320 of 8000 = 320/8000 *100
= 4%